OpenStudy (anonymous):

Clear the fractions to solve. 1/3x+ 5/6x+ 4/3= 7/6x= 10/3

OpenStudy (luigi0210):

Multiply by a common factor.

OpenStudy (luigi0210):

What do they all have in common?

OpenStudy (anonymous):

would that be 3?

OpenStudy (luigi0210):

If we multiplied by 3 we would still have fractions

OpenStudy (anonymous):

ok.... what does that mean? like 3/1? or something?

OpenStudy (luigi0210):

Multiply by a number that would cancel out the 3 and 6.

OpenStudy (luigi0210):

Question, why does the question have 2 equal signs?

OpenStudy (anonymous):

it was suppose to be a + sign for the first one the second one is right

OpenStudy (luigi0210):

Oh alright. But yea the best way to clear the fractions would be to multiply by 6.

OpenStudy (anonymous):

ok and after that? @luigi0210

OpenStudy (luigi0210):

Just get variables to one side and numbers to the other

OpenStudy (anonymous):

U multiply by largest common division factor to clear fraction. this case 6 so (1/3x+ 5/6x+ 4/3= 7/6x= 10/3) x 6 <- multiply 6 both side. u r done.